'Bhartdarshan’ is an Internet-based travel agency wherein customers can see videos of the


cities they plan to visit. The number of hits daily is a normally distributed random variable


with a mean of 10,000 and a standard deviation of 2,400.


a. What is the probability of getting more than 12,000 hits?



1
Expert's answer
2022-04-17T23:28:50-0400

"P(X>12000)=P(Z>\frac{12000-10000}{2400})=P(Z>0.83)=1-P(Z<0.83)=0.2033."